ADNAN MUHAMMED

Grade 12,

One card is drawn from a well-shuffled pack of 52 cards. What is the probability that a card will be (i) a diamond (ii) Not an ace (iii)a black card (iv) not a diamond

One card is drawn from a well-shuffled pack of 52 cards. What is the probability that a card will be

(i) a diamond (ii) Not an ace (iii)a black card (iv) not a diamond

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

i) the probability that a card is a diamond

We know that there are 13 diamond cards in a deck.
Therefore, the required probability is:

P( getting a diamond card) = 13/52
= ¼

(ii) the probability that a card is not an ace

We know that there are 4 ace cards in a deck.

Therefore, the required probability is:

P(not getting an ace card) = 1-( 4/52)

= 1- (1/13)

=(13-1)/13

= 12/13

(iii) the probability that a card is a black card

We know that there are26 black cards in a deck.

Therefore, the required probability is:

P(getting a black card) = 26/52 = ½

(iv) the probability that a card is not a diamond

We know that there are 13 diamond cards in a deck.

We know that the probability of getting a diamond card is 1/4

Therefore, the required probability is:

P( not getting a diamond card) = 1- (1/4)

= (4-1)/4

= ¾

Thanks